# TokenWeave: Efficient Compute-Communication Overlap for Distributed LLM Inference

### v1

@@
Distributed inference of large language models (LLMs) can introduce overheads of up to 20% even over GPUs connected via high-speed interconnects such as NVLink. Multiple techniques have been proposed to mitigate these overheads by decomposing computations into finer-grained tasks and overlapping communication with subtasks as they complete. However, fine-grained decomposition of a large computation into many smaller computations on GPUs results in overheads. Further, the communication itself uses many streaming multiprocessors (SMs), adding to the overhead.
We present TokenWeave to address these challenges. TokenWeave proposes a Token-Splitting technique that divides the tokens in the inference batch into two approximately equal subsets in a wave-aware manner. The computation of one subset is then overlapped with the communication of the other. In addition, TokenWeave optimizes the order of the layer normalization computation with respect to communication operations and implements a novel fused AllReduce-RMSNorm kernel that carefully leverages Multimem instruction support available on NVIDIA Hopper GPUs. These optimizations allow TokenWeave to perform communication and RMSNorm using only 2-8 SMs. Moreover, our kernel enables the memory bound RMSNorm to be overlapped with the other batch’s computation, providing additional gains. Our evaluations demonstrate up to 29% latency gains and up to 26% throughput gains across multiple models and workloads. In several settings, TokenWeave results in better performance compared to an equivalent model with all communication removed. The source code of TokenWeave will be released soon.
@@ 1. Introduction
Figure 1. Communication overhead of AllReduce for three models vs sequence length on 8xH100 DGX (error bar shows std. dev. across 5 runs). Even with NVLink/ NVSHARP, communication overheads can be over 20%.
Figure 2. Inference latency of Llama-3.3-70B on 8xH100 DGX for various sequence lengths. vllm-Multimem corresponds to vLLM with an optimized AllReduce implementation using Multimem and NVSHARP support. vllm-nocomm is a counterfactual baseline corresponding to only the computation time without any communication. The dotted lines represent relative performance compared to the vllm-Multimem baseline with TokenWeave providing up to 29% gains. Even at smaller sequence lengths, TokenWeave provides significant gains, e.g. $18\%$ at $1K$, while prior schemes incur overheads. At sequence lengths $\geq 4K$, TokenWeave outperforms vllm-nocomm by not only recovering the complete communication overhead but also providing additional gains due to our fused AllReduce-RMSNorm kernel.
Modern LLM deployments increasingly rely on distributed inference across multiple GPUs, driven by model size constraints or stringent latency requirements. Although high-speed interconnects like NVLink significantly improve communication bandwidth, substantial overheads remain. For example, in our experiments communication accounted for up to 23% of the end-to-end inference latency for Llama-3.3-70B (Grattafiori et al., 2024) and Qwen2.5-72B (Yang et al., 2024) models running on 8xH100 DGX (Figure 1).
@@
In this paper, we present TokenWeave, an efficient method to minimize communication overhead for LLM inference. To the best of our knowledge, TokenWeave is the first system that can significantly reduce communication overheads even in batches with moderate number of tokens. For example, as shown in Figure 2 with $tokens\_in\_batch=1024$, TokenWeave achieves latency improvements of $18\%$ on Llama-3.3-70B on 8xH100. On the other hand, state-of-the-art communication fusion methods like TileLink (Zheng et al., 2025), for the same setting, results in an overall hit in performance of 50+over the non-overlapped baseline (Figure 14). Reducing the $tokens\_in\_batch$ requirement makes TokenWeave a viable technique to use together with modern schedulers such as Sarathi (Agrawal et al., 2024) which uses chunked prefill for efficient inference. For example, the popular vLLM (Kwon et al., 2023) serving system uses a default chunked prefill size (i.e. max_num_batched_tokens) of 2048 (vLLM, [n. d.]). Even in disaggregated settings like Splitwise (Patel et al., 2024) or DistServ (Zhong et al., 2024) with $tokens\_in\_batch\geq 4096$, Figure 2 shows that TokenWeave delivers over 20% gains and outperforms an equivalent model that has all communication removed.
@@
TokenWeave comprises of three key techniques. First, we propose Token-Splitting, which is a coarse-grained splitting of tokens in the input batch into just two approximately equal subsets so that the computation of one split can be overlapped with communication of the other. The token-based splitting (instead of request-based) allows us to split in either the batch- or sequence- dimensions independently. Naïvely splitting the input tokens into two equal halves can result in overheads due to wave-quantization. Thus, TokenWeave performs Smart-splitting, a GPU wave-aware splitting which ensures that the total number of waves executed in kernels of the two subtasks is not more than the number of waves in the full computation kernel. The coarse work partitioning and smart-splitting almost completely eliminates the compute overheads due to the smaller computations.
@@
Second, while state-of-the-art models today use the more efficient RMSNorm (Zhang and Sennrich, 2019) instead of LayerNorm for normalization, we find that even RMSNorm can take as much as 8% of the overall latency on H100s (Figure 3). Thus, unlike previous work such as Tilelink (Zheng et al., 2025), Flux (Chang et al., 2024) and Nanoflow (Zhu et al., 2024) that does not optimize normalization operations, TokenWeave carefully optimizes the normalization computation along with the communication operation.
@@
Figure 3. RMSNorm overhead for different models vs sequence length on 8xH100 DGX (error bar shows std. dev. across 5 runs). RMSNorm performed after AllReduce has non-trivial overheads and can range from $5-9\%$.
@@
Normalization operations are token-level operations that normalize the distribution of values along the hidden dimension and are memory-bandwidth bound. By default, on LLM serving systems like vLLM, normalization is performed after the AllReduce communication operation in most models. Logically, AllReduce can be broken into (equivalent) ReduceScatter and AllGather operations. If RMSNorm is moved in between these two operations, it does not change the mathematical function, but reduces the RMSNorm’s computational requirement by a factor of the number of GPUs in the distributed replica. However, simply breaking AllReduce into ReduceScatter and AllGather results in significant overhead that annuls the gains of the reduced RMSNorm computation, resulting in no advantage of doing such a breakup (Figure 4). In TokenWeave, we design an optimized kernel that fuses communication and normalization such that it utilizes minimal number of HBM reads and writes, resulting in a very efficient implementation.
@@
Third, we leverage modern hardware features to implement communication with minimal SM use. For ReduceScatter, we leverage the Multimem instruction which offloads the reduction operation to the NVSwitch, reducing the SM requirement. For AllGather, we rely on the Multimem store instructions which reduces both memory-bandwidth and communication cost, minimizing SM requirement again. Due to the reduced computation and memory bandwidth requirements, the AllReduce-RMSNorm fused kernel can be efficiently executed with just 2-8 SMs compared to 16-20 SMs used by communication in prior work (Zhu et al., 2024). Figure 4 shows up to 40% gains due to this fusion across the range of sequence lengths. An additional benefit of the fused kernel is that the memory-bandwidth bound RMSNorm operation can now also be overlapped with the computation of the other batch.
@@
Figure 4. Latency of executing a single AllReduce and RMSNorm operation for 8192 dimension hidden size, bf16 on 8xH100 DGX across varying sequence lengths for three approaches. Simple reordering of RMSNorm between ReduceScatter (RS) and AllGather (AG) is worse than performing RMSNorm after AllReduce except at high sequence lengths as the overheads of splitting AllReduce (AR) into RS and AG eat into the gains. Our fused AllReduce-RMSNorm kernel results in up to 40% gains over the entire sequence length range.
@@
We have integrated TokenWeave into the latest vLLM V1 serving framework and evaluated its performance across various models. We show that TokenWeave achieves up to 29% gains in overall latency compared to an optimized non-overlapped baseline across a range of batch sizes. In comparison, state-of-the-art communication overlap solution, TileLink, results in latency improvements only for batch sizes $>=4K$ and that too with much more modest gains. Further, for various end-to-end workloads, we find that TokenWeave delivers up to 26% gains in throughput. Finally, TokenWeave is able to outperform an equivalent model with all communication removed due to its novel fused AllReduce-RMSNorm kernel that optimizes and overlaps the memory-bandwidth bound normalization operation.
@@
    A novel fused AllReduce-RMSNorm kernel that leverages modern GPU hardware capabilities to jointly optimize layer normalization and communication, thereby requiring only 2-8 SMs.
@@
    Implementation on the latest vLLM V1 inference system and extensive evaluation over multiple models, demonstrating improvements of up to 29% latency gains and up to 26% throughput gains.
@@ 2.1. Distributed Inference
TP thus requires two AllReduce operations per transformer block which lie in the critical path. This can add significant cost to inference latency and reduce GPU efficiency. For example, as shown in Figure 1, the communication cost can add an overhead of up to $23\%$ even on an H100 DGX which have high-speed NVLink interconnects.
@@ 2.2. Compute-Communication Overlap
Figure 5. Splitting AllReduce (AR) into ReduceScatter (RS) and AllGather (AG) can result in non-trivial overheads. Shown are the individual times and the relative performance (line plots) of these operations on 8xH100 DGX. All runs are with a hidden size of 8192 with bf16.
@@
Figure 6. Large collective operations are more efficient. Shown is the bandwidth for ReduceScatter (RS) on 8xH100 DGX for varying sequence length (hidden size 8192, bf16). Larger tensors result in much better bandwidth, demonstrating that splitting input into smaller parts results in overheads.
@@
The fused kernel approaches (Wang et al., 2022; Chang et al., 2024; Zheng et al., 2025; Hassani et al., [n. d.]; Wang et al., [n. d.]), suffer from three problems. First, since the techniques rely on overlap within a GEMM kernel, communication overlap during non-GEMM operations like attention is not feasible. For example, the AllReduce after FFN layer is performed by overlapping the ReduceScatter portion with the second MLP, and overlapping the AllGather portion with the QKV-preprojection of the next attention layer. This limits the overlap opportunity and for small models or for small batches where the QKV-projection takes a very small time, there is not enough time to overlap the AllGather. Second, the splitting into ReduceScatter and AllGather operations, is less efficient compared to the equivalent AllReduce operation. Figure 5 plots the performance of a single AllReduce kernel compared to the cost of the equivalent ReduceScatter+AllGather for different sizes of communication. As shown, this can add significant overheads, upwards of 50%. Third, the smaller tile-sized granularity of communication is less efficient than doing single large transfers. Figure 6 plots the achieved ReduceScatter bandwidth vs the tensor size. As shown, smaller communication sizes achieve much lower bandwidths. Due to these factors,  (Wang et al., 2022; Chang et al., 2024; Zheng et al., 2025) are able to effectively hide communication only when the GEMM is large enough, requiring batches with large tokens-in-batch.
@@ 2.3. NV-SHARP, Multicast and SymmetricMemory
These hardware and software advancements considerably decrease the number of SMs required for executing communication primitives and alleviate memory bandwidth pressure. In our experiments, we observed that utilizing only 6-8% of SMs on the H100 GPUs suffices to saturate communication bandwidth, leaving the majority of SMs free for compute tasks that overlap communication. This can be seen in Figure 7 which shows the latency of AllReduce kernel vs the number of SMs.
@@ 3. TokenWeave: Design and Implementation
Figure 8. Overview of TokenWeave. (a) Vanilla Tensor Parallelism: all compute and communication operations are performed sequentially. (b) TokenWeave: input batch is partitioned into two splits. RMSNorm is fused with communication and communication of one split overlaps with computation of the other split. Separate compute and communication streams blend to orchestrate the overlap.
@@
We now describe the main techniques of TokenWeave– a coarse grained smart Token-Splitting technique, RMSNorm reordering and a fused AllReduce-RMSNorm implementation. Figure 8(b) provides a high-level schematic of our method compared to the standard tensor parallel implementation.
@@ 3.1. Coarse-Grained Token-Splitting
TokenWeave partitions the incoming batch into two subsets, each with nearly equal computational and communication requirement. For illustration, consider an incoming batch with size 1 and a sequence comprising $L$ tokens. This batch is divided into two split-batches: a prefix-split containing the initial subsequence of length $L_{1}$, and a suffix-split containing the remaining subsequence of length $L_{2}$ ($L=L_{1}$ + $L_{2}$). These split-batches are processed separately in a pipelined manner.
@@ 3.1.1. Wave-Aware Smart-splitting
Partitioning a large computation into smaller units can introduce overhead on GPUs due to wave quantization effects. Consider a GEMM kernel that requires 300 CTAs (Cooperative Thread Arrays). On an NVIDIA H100 GPU with 132 SMs (Streaming Multiprocessors), assuming each CTA occupies exactly one SM, this computation would span two full waves and one partial wave utilizing 32 SMs. Thus, the total computation time equals three waves of execution.

### v2

@@
Distributed inference of large language models (LLMs) can introduce overheads of up to $20\%$ even over GPUs connected via high-speed interconnects such as NVLink. Multiple techniques have been proposed to mitigate these overheads by decomposing computations into finer-grained tasks and overlapping communication with subtasks as they complete. However, fine-grained decomposition of a large computation into many smaller computations on GPUs results in overheads. Furthermore, the communication itself uses many streaming multiprocessors (SMs), adding to the overhead.
We present TokenWeave to address these challenges. TokenWeave proposes a Token-Splitting technique that divides the tokens in the inference batch into two approximately equal subsets in a wave-aware manner. The communication of one subset is then overlapped with the computation of the other. In addition, TokenWeave optimizes the order of the layer normalization computation with respect to communication operations and implements a novel fused AllReduce–RMSNorm [^1] In all modern implementations, residual addition is fused with RMSNorm, and all our results use this fused implementation. For brevity, we mention only RMSNorm in all our discussions. kernel that carefully leverages Multimem instruction support available on NVIDIA Hopper GPUs. These optimizations allow TokenWeave to perform communication and RMSNorm using only $2-8$ SMs. Moreover, our kernel enables the memory-bound RMSNorm to be overlapped with the other batch’s computation, providing additional gains.
Our evaluations demonstrate up to $\boldsymbol{1.29\times}$ speedup in latency and $\boldsymbol{1.26\times}$ higher throughput across multiple models and workloads. In several settings, TokenWeave results in better performance compared to an equivalent model with all communication removed. The source code of TokenWeave is available at https://github.com/microsoft/tokenweave.
@@ 1. Introduction
Figure 1. Communication overhead of AllReduce for three models vs. sequence length on an 8×8$\times$8 ×H100 DGX (error bar shows std. dev. across 5 runs). Even with NVLink/NVSHARP, communication overheads can be over $20\%$.
Figure 2. Inference latency of Llama-3.3-70B on 8xH100 DGX for various sequence lengths. vllm-Multimem corresponds to vLLM with an optimized AllReduce implementation using Multimem and NVSHARP support. vllm-nocomm is a counterfactual baseline corresponding to only the computation time without any communication. The dotted lines show performance normalized to the vllm-Multimem baseline. TokenWeave achieves up to 1.29×1.29$\times$1.29 × speedup. Even at shorter sequence lengths, TokenWeave provides significant gains, for example, 1.18×1.18$\times$1.18 × at a sequence length of $1K$ tokens, while prior schemes incur overheads. At sequence lengths $\geq 4K$, TokenWeave outperforms vllm-nocomm by not only recovering the complete communication overhead but also providing additional gains due to our fused AllReduce–RMSNorm kernel.
Modern LLM deployments increasingly rely on distributed inference across multiple GPUs, driven by model size constraints or stringent latency requirements. Although high-speed interconnects like NVLink significantly improve communication bandwidth, substantial overheads remain. For example, in our experiments communication accounted for up to $23\%$ of the end-to-end inference latency for Llama-3.3-70B (Grattafiori et al., 2024) and Qwen2.5-72B (Yang et al., 2024) models running on 8×8$\times$8 ×H100 DGX (Figure 1).
@@
In this paper, we present TokenWeave, an efficient method to minimize communication overhead for LLM inference. To the best of our knowledge, TokenWeave is the first system that can significantly reduce communication overheads even in batches with moderate number of tokens. For example, as shown in Figure 2 with $tokens\_in\_batch=1024$, TokenWeave achieves a 1.18×1.18$\times$1.18 × latency improvement on Llama-3.3-70B running on 8×8$\times$8 ×H100 GPUs. In contrast, state-of-the-art communication fusion methods such as TileLink (Zheng et al., 2025), under the same setting, achieve only 0.51×0.51$\times$0.51 × of the non-overlapped baseline performance, representing a significant degradation (Figure 14). Reducing the $tokens\_in\_batch$ requirement makes TokenWeave a viable technique to use together with modern schedulers such as Sarathi (Agrawal et al., 2024) which uses chunked prefill for efficient inference. For example, the popular vLLM (Kwon et al., 2023) serving system uses a default chunked prefill size (i.e. max_num_batched_tokens) of $2048$ (vLLM, [n. d.]). Even in disaggregated settings like Splitwise (Patel et al., 2024) or DistServe (Zhong et al., 2024) with $tokens\_in\_batch\geq 4096$, Figure 2 shows that TokenWeave delivers >1.20×>1.20$\times$> 1.20 × improvement and outperforms an equivalent model that has all communication removed.
@@
TokenWeave comprises three key techniques. First, we propose Token-Splitting, which is a coarse-grained splitting of tokens in the input batch into just two approximately equal subsets so that the computation of one split can be overlapped with the communication of the other. The token-based splitting (instead of request-based) allows us to split in either the batch- or sequence- dimensions independently. Naïvely splitting the input tokens into two equal halves can result in overheads due to wave-quantization. Thus, TokenWeave performs Smart-splitting, a GPU wave-aware splitting which ensures that the total number of waves executed in kernels of the two subtasks is not more than the number of waves in the full computation kernel. The coarse work partitioning and smart-splitting almost completely eliminate the compute overheads due to the smaller computations.
@@
Second, while state-of-the-art models today use the more efficient RMSNorm (Zhang and Sennrich, 2019) instead of LayerNorm for normalization, we find that even RMSNorm can take as much as $8\%$ of the overall latency on H100s (Figure 3). Thus, unlike previous work such as Tilelink (Zheng et al., 2025), Flux (Chang et al., 2024) and Nanoflow (Zhu et al., 2024) that do not optimize normalization operations, TokenWeave carefully optimizes the normalization computation along with the communication operation.
@@
Figure 3. RMSNorm overhead for different models vs sequence length on an 8×8$\times$8 ×H100 DGX (error bar shows std. dev. across 5 runs). RMSNorm performed after AllReduce has non-trivial overheads and can range from $5-9\%$.
@@
Normalization operations are token-level operations that normalize the distribution of values along the hidden dimension and are memory-bandwidth bound. By default, on LLM serving systems like vLLM, normalization is performed after the AllReduce communication operation in most models. Logically, AllReduce can be broken into (equivalent) ReduceScatter and AllGather operations. If RMSNorm is moved in between these two operations, it does not change the mathematical function, but reduces the RMSNorm’s computational requirement by a factor of the number of GPUs in the distributed replica. However, simply breaking AllReduce into ReduceScatter and AllGather results in significant overhead that annuls the gains of the reduced RMSNorm computation, resulting in no advantage of doing such a breakup (Figure 4). In TokenWeave, we design an optimized kernel that fuses communication and normalization such that it utilizes a minimal number of HBM reads and writes, resulting in a very efficient implementation.
@@
Third, we leverage modern hardware features to implement communication with minimal SM use. For ReduceScatter, we leverage the Multimem instruction which offloads the reduction operation to the NVSwitch, reducing the SM requirement. For AllGather, we rely on the Multimem store instructions which reduce both memory bandwidth and communication cost, minimizing SM requirement again. Due to the reduced computation and memory bandwidth requirements, the AllReduce–RMSNorm fused kernel can be efficiently executed with just $2-8$ SMs compared to $16-20$ SMs used by communication in prior work (Zhu et al., 2024). Figure 4 shows up to 1.40×1.40$\times$1.40 × improvement due to this fusion across the range of sequence lengths. An additional benefit of the fused kernel is that the memory bandwidth bound RMSNorm operation can now also be overlapped with the computation of the other batch.
@@
Figure 4. Latency of executing a single AllReduce and RMSNorm operation for $8192$ dimension hidden size, $bf16$ on 8×8$\times$8 ×H100 DGX across varying sequence lengths for three approaches. Simple reordering of RMSNorm between ReduceScatter (RS) and AllGather (AG) is worse than performing RMSNorm after AllReduce except at high sequence lengths as the overheads of splitting AllReduce (AR) into RS and AG eat into the gains. Our fused AllReduce–RMSNorm kernel results in up to 1.40×1.40$\times$1.40 × improvement over the entire sequence length range.
@@
We have integrated TokenWeave into the latest vLLM-V1 serving framework and evaluated its performance across various models. We show that TokenWeave achieves up to 1.29×1.29$\times$1.29 × better overall latency compared to an optimized non-overlapped baseline across a range of batch sizes. In comparison, state-of-the-art communication overlap solution, TileLink, results in latency improvements only for batch sizes $>=4K$ and that too with much more modest gains. Further, for various end-to-end workloads, we find that TokenWeave delivers up to 1.26×1.26$\times$1.26 × higher throughput. Finally, TokenWeave is able to outperform an equivalent model with all communication removed due to its novel fused AllReduce–RMSNorm kernel that optimizes and overlaps the memory-bandwidth bound normalization operation.
@@
    A novel fused AllReduce–RMSNorm kernel that leverages modern GPU hardware capabilities to jointly optimize layer normalization and communication, thereby requiring only $2-8$ SMs.
@@
    Implementation on the latest vLLM-V1 inference system and extensive evaluation across multiple models demonstrate up to 1.29×1.29$\times$1.29 × latency improvement and up to 1.26×1.26$\times$1.26 × throughput improvement.
@@ 2.1. Distributed Inference
TP thus requires two AllReduce operations per transformer block which lie in the critical path. This can add significant cost to inference latency and reduce GPU efficiency. For example, as shown in Figure 1, the communication cost can add an overhead of up to $23\%$ even on an 8×8$\times$8 ×H100 DGX which has high-speed NVLink interconnects.
@@ 2.2. Compute-Communication Overlap
Figure 5. Splitting AllReduce (AR) into ReduceScatter (RS) and AllGather (AG) can result in non-trivial overheads. Shown are the individual times and the relative performance (line plots) of these operations on 8×8$\times$8 ×H100 DGX. All runs are with a hidden size of $8192$ with $bf16$.
@@
Figure 6. Large collective operations are more efficient. Shown is the bandwidth for ReduceScatter (RS) on 8×8$\times$8 ×H100 DGX for varying sequence length (hidden size $8192$, $bf16$). Larger tensors result in much better bandwidth, demonstrating that splitting input into smaller parts results in overheads.
@@
The fused kernel approaches (Wang et al., 2022; Chang et al., 2024; Zheng et al., 2025; Hassani et al., [n. d.]; Wang et al., [n. d.]), suffer from three problems. First, since the techniques rely on overlap within a GEMM kernel, communication overlap during non-GEMM operations like attention is not feasible. For example, the AllReduce after FFN layer is performed by overlapping the ReduceScatter portion with the second MLP, and overlapping the AllGather portion with the QKV-preprojection of the next attention layer. This limits the overlap opportunity and for small models or for small batches where the QKV-projection takes a very small time, there is not enough time to overlap the AllGather. Second, the splitting into ReduceScatter and AllGather operations, is less efficient compared to the equivalent AllReduce operation. Figure 5 plots the performance of a single AllReduce kernel compared to the cost of the equivalent ReduceScatter+AllGather for different sizes of communication. As shown, this can add significant overheads, upwards of $50\%$. Third, the smaller tile-sized granularity of communication is less efficient than doing single large transfers. Figure 6 plots the achieved ReduceScatter bandwidth vs the tensor size. As shown, smaller communication sizes achieve much lower bandwidths. Due to these factors,  (Wang et al., 2022; Chang et al., 2024; Zheng et al., 2025) are able to effectively hide communication only when the GEMM is large enough, requiring batches with large tokens-in-batch.
@@ 2.3. NV-SHARP, Multicast and SymmetricMemory
These hardware and software advancements considerably decrease the number of SMs required for executing communication primitives and alleviate memory bandwidth pressure. In our experiments, we observed that utilizing only $6-8\%$ of SMs on the H100 GPUs suffices to saturate communication bandwidth, leaving the majority of SMs free for compute tasks that overlap communication. This can be seen in Figure 7 which shows the latency of AllReduce kernel vs the number of SMs.
@@ 3. TokenWeave: Design and Implementation
Figure 8. Overview of TokenWeave. (a) Vanilla Tensor Parallelism: all compute and communication operations are performed sequentially. (b) TokenWeave: input batch is partitioned into two splits. RMSNorm is fused with communication and communication of one split overlaps with computation of the other split. Separate compute and communication streams weave to orchestrate the overlap.
@@
We now describe the main techniques of TokenWeave—a coarse-grained, smart Token-Splitting technique, Resi-
dualAdd–RMSNorm reordering, and a fused AllReduce–RMSNorm implementation. Figure 8(b) provides a high-level schematic of our method compared to the standard tensor-parallel implementation.
@@ 3.1. Coarse-Grained Token-Splitting
TokenWeave partitions the incoming batch into two subsets, each with nearly equal computational and communication requirements. For illustration, consider an incoming batch with size 1 and a sequence comprising $L$ tokens. This batch is divided into two split-batches: a prefix-split containing the initial subsequence of length $L_{1}$, and a suffix-split containing the remaining subsequence of length $L_{2}$ ($L=L_{1}$ + $L_{2}$). These split-batches are processed separately in a pipelined manner.
@@ 3.1.1. Wave-Aware Smart-splitting
Partitioning a large computation into smaller units can introduce overhead on GPUs due to wave quantization effects. Consider a GEMM kernel that requires 300 CTAs (Cooperative Thread Arrays). On an NVIDIA H100 GPU with 132 SMs (Streaming Multiprocessors), assuming each CTA occupies exactly one SM, this computation would span two full waves and one partial wave utilizing 36 SMs. Thus, the total computation time equals three waves of execution.